xxcoffee:

What is the slope of the line 3y = 6x - 1? a. -1 b. 1/2 c. 2 d. 3

yesno:

divide both sides by 3. y=2x-(1/3) Slope is the number in the position of "m", in linear equation y=mx+b. Slope= 2. or the answer would be... 2.

xxcoffee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @yesno divide both sides by 3. y=2x-(1/3) Slope is the number in the position of "m", in linear equation y=mx+b. Slope= 2. or the answer would be... 2. \(\color{#0cbb34}{\text{End of Quote}}\) ight thxx i got that one now

yesno:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxcoffee \(\color{#0cbb34}{\text{Originally Posted by}}\) @yesno divide both sides by 3. y=2x-(1/3) Slope is the number in the position of "m", in linear equation y=mx+b. Slope= 2. or the answer would be... 2. \(\color{#0cbb34}{\text{End of Quote}}\) ight thxx i got that one now \(\color{#0cbb34}{\text{End of Quote}}\) mhm yea no problem.